# arXiv:2603.01660v1[eess.SP] 02 Mar 2026

### v1

@@ Abstract
Non-Line-of-Sight (NLoS) sensing and detection of low-observable (stealth) targets are challenging for conventional radar due to blockage and severe propagation loss. Intelligent Reflective Surface (IRS)-assisted radar can extend the field-of-view (FOV), but common architectures rely on the four-hop radar–IRS–target–IRS–radar link, whose attenuation limits estimation performance. This paper proposes an alternative architecture, that exploits the target-scattered component received at a spatially separated IRS and redirected back to a mono-static radar receiver. The geometry provides bi-static/multi-static-like diversity using a passive panel, while retaining a mono-static front-end and avoiding inter-node time synchronization concerns. We develop a signal model for the proposed configuration and recast it into a compact, parameterized form that is suitable for angle estimation. Using this reformulation, we derive the Fisher Information Matrix and the associated Cramér–Rao Lower Bounds (CRLB) for target azimuth and elevation angles with respect to the IRS. Numerical evaluations quantify the impact of various signal-model parameters on the achievable bounds. These results provide insights on the parameter-estimation limits within the FOV against SNR, snapshots and IRS elements.
@@ I Introduction
Radar has historically been a vital sensing technology capable of active and passive detection, target tracking, high-res imaging, and increasingly, sensing in Non-Line-of-Sight (NLoS) environments[16, 13, 22, 18]. Conventional mono-static radars usually perform well when targets are within their Line-of-Sight (LoS), enabling precise measurement of range, Doppler shifts, and directions. However, when obstacles block the LoS, the effectiveness of traditional radar diminishes significantly or results in complete detection failure, driving the innovation of new NLoS sensing architectures.
A common method for enabling NLoS radar sensing involves using an Intelligent Reflective Surface (IRS). An IRS is made up of passive, adaptable reflecting elements that can steer incident electromagnetic waves, creating controllable pathways beyond the LoS when positioned appropriately. This IRS-assisted radar technology can be well suited in defence, automotive sensing, and modern wireless systems to expand the field-of-view (FOV) and improve NLoS target detection capabilities [21, 19, 4]. Meanwhile, the presence of an IRS introduces multiple possible signal paths from the target to the radar. Specifically, deploying an IRS results in four key propagation paths (which vary based on geometry and blockage), that include:
@@
Most IRS-assisted radar studies typically focus on the dominant LoS component (P1) and the IRS-enabled NLoS component (P2) [5]. NLoS sensing via (P2) has been widely studied, including the derivation of Cramér–Rao Lower Bounds (CRLB) for selected target parameters [14]. However, while IRS phase weights can provide beamforming gain, they cannot fundamentally mitigate the severe path loss inherent to the four-hop channel in (P2), due to compounded attenuation over the radar–IRS and IRS–target links [12].
@@
To mitigate performance degradation due to path loss and scattering from stealth targets, radar system design has also inclined towards bi-static and multi-static configurations. Unlike mono-static radar, in which the transmitter (Tx) and receiver (Rx) are co-located, bi-static and multi-static radars spatially separate the Tx and Rx(s), thereby forming a distributed sensing architecture [10, 17, 6]. This separation enables the receiver to observe the target from a viewpoint that is not constrained by the LoS path between the transmitter and target. In multi-static reception, multiple spatially separated receivers observe the target returns, which can improve detection probability and tracking performance through spatial diversity [20, 1]. Such configurations are particularly useful for detecting low-observable (stealth) targets and for sensing space objects, making them attractive for modern defence and space applications [15, 7]. Despite these advantages, they introduce important practical limitations, most notably stringent time and phase synchronization requirements between the Tx and distributed Rx (s), along with increased hardware cost and maintenance overhead compared to mono-static radar.
@@
Upon investigating these trade-offs, we observe the need for a configuration that (i) reduces the effective path loss relative to the double-IRS (P2) route, and (ii) avoids inter-node time synchronization errors inherent in bi-static/multi-static radars. In this work, we focus on the third path (P3), i.e., Radar $\rightarrow$ Target $\rightarrow$ IRS $\rightarrow$ Radar, which is often overlooked but is extremely useful. As will be shown, this path is analogous to a bi-static/multi-static observation geometry because the IRS provides a spatially displaced “listening” viewpoint relative to the radar, yet it remains a passive structure. This motivates a bi-static-like configuration based on IRS-assisted radar. A similar configuration is presented in  [3] for stealth target detection. However, parameter estimation in such a scenario is not discussed.
@@
In the proposed IRS-assisted Radar configuration from  [3], an IRS is positioned spatially away from the mono-static radar. The radar transmits a signal into its FOV; the target reflects and scatters the incident energy. The IRS is oriented (and configured) to “look” toward a specific direction that intersects the radar’s FOV at a different angle, receive the scattered signal from the target, and steer it back to the radar receiver. In this manner, the sensing chain leverages (P3), mimicking certain advantages of bi-static/multi-static radar while maintaining a mono-static radar front-end, thereby eliminating explicit time synchronization errors between Tx and Rx. Moreover, employing an IRS panel to realize a bi-static-like sensing geometry in challenging scenarios (e.g., stealth target detection) introduces a novel paradigm: from a geometric sensing standpoint, the IRS-assisted propagation path can often be interpreted as effectively translating the sensing aperture to the IRS location and subsequently mapping the estimated target coordinates back into the original radar reference frame.
@@
  • The IRS-assisted radar configuration is presented in comparison to bi-static/multi-static radar and existing IRS-assisted radars. The associated signal model is systematically developed and reformulated into a parameterized model suitable for CRB/CRLB derivation.
  • We derive Cramér–Rao bounds (CRB) for the target angle parameters (azimuth and elevation) under the proposed signal model.
  • Extensive simulation studies are performed to validate the performance of the radar system under different geometric and channel conditions.
The paper is organized as follows: Section II explains in detail the proposed radar configuration. Section III provides the detailed mathematical derivation of the signal model and the target parameters. Based on these target parameters, the theoretical CRB limits are derived in Section IV. Section V discusses the numerical results in relation to the theoretical bounds, and Section VI concludes the paper.
@@ II IRS-Assisted Radar Setup
Bi-static radar separates the transmitter and receiver, forming a triangle with the target, unlike mono-static radar’s single-axis setup. This concept generalizes to multi-static radar, which uses multiple spatially dispersed transmitters and receivers to gain spatial diversity. Multi-static systems observe targets from several angles, enhancing the detection of complex or stealthy targets whose Radar Cross Section (RCS) varies with aspect. However, effectively exploiting these advantages requires careful spatial arrangement of the receivers, as discussed in [9].
The IRS-assisted radar setup replaces traditional bi-/multi-static receiver nodes with low-cost IRS panels that reflect scattered signals toward nearby radar sensors  [3]. By carefully designing the IRS element weights, the system achieves sufficient beamforming gain to offset longer propagation paths. Fixed IRS placement simplifies radar-IRS channel estimation, while the large number of IRS elements enhances parameter estimation and mitigates synchronization issues typical in conventional bi-/multi-static radars, which rely on fewer, more expensive antennas.
The text introduces an “IRS-static radar” configuration, based on the IRS-assisted radar setup in [2], that uses only two relevant paths: the monostatic radar path and the Radar → Target → IRS → Radar path, with the latter being crucial for stealth targets. Unlike prior IRS-assisted radar work [8], which places the IRS near the radar or target to boost signal strength and relies on a 4-path model (Radar → IRS → Target → IRS → Radar)[cite 4 path model papers], the proposed design positions the IRS panel away from the radar and operates it in simplex mode so that energy always flows Target → IRS → Radar. This yields a simpler IRS (lower design and circuitry complexity), forms an effective 3-path model (Radar → Target → IRS → Radar), reduces path loss compared to conventional 4-path IRS-assisted systems, and functionally emulates a bi-static radar configuration, hence the term “IRS-static radar.”
@@
Figure 1: Comparison of radar configurations (Left) and the IRS-static radar setup (Right).
@@ III Signal Model
Consider an air-target scenario with $K$ targets defined with respect to the radar coordinate system, with the radar located at $\mathrm{r}_{\mathrm{R}}=[0,0,0]^{T}$. Let the $k$th target parameters—namely position, velocity, azimuth angle, and elevation angle—be given by $\mathbf{r}_{k}=[x_{k},y_{k},z_{k}]^{T}$, $\mathbf{v}_{k}=[v_{kx},v_{ky},v_{kz}]^{T}$, $\theta^{(\mathrm{az})}_{k}$, and $\theta^{(\mathrm{el})}_{k}$, respectively. Without loss of generality, we consider a single IRS panel with $L_{h}\times L_{v}$ reflecting elements located at $\mathbf{r}_{\mathrm{I}}$ in the radar coordinate system. Let the $k$th target subtend azimuth and elevation angles $\psi^{(\mathrm{az})}_{k}$ and $\psi^{(\mathrm{el})}_{k}$, respectively, as seen from the IRS coordinate system, as illustrated in Fig. LABEL:system. We assume a phased-array radar with $M_{T}=M_{R}$ transceivers co-located at $\mathbf{r}_{\mathrm{R}}$.
@@
Next, we define the line-of-sight (LoS) channel between the $k$th target and the IRS panel as $\mathbf{h}_{k}\in\mathbb{C}^{L_{h}L_{v}\times 1}$, and the channel between the IRS panel and the phased-array transceiver as $\mathbf{F}\in\mathbb{C}^{M_{R}\times L}$. For a 2-D IRS panel, $L=L_{h}L_{v}$, where $L_{h}$ and $L_{v}$ denote the number of elements along the horizontal and vertical dimensions, respectively. The signal received at the radar via the IRS path (denoted as “is”) is given by
$\mathbf{y}^{\mathrm{is}}(t,t_{l})=\sum_{k=1}^{K}\alpha^{\mathrm{is}}_{k}\mathbf{F}\mathbf{D}\mathbf{h}_{k}{\mathbf{g}_{k}}^{H}\mathbf{s}(t-\tau_{k,l}),$ (2)
where $\alpha^{\mathrm{is}}_{k}$ is the target response strength through the IRS path and $\mathbf{D}=\mathrm{diag}\left(d_{1},\dots,{d}_{l},\dots,{d}_{L_{h}L_{v}}\right)\ \in\mathbb{C}^{L_{h}L_{v}\times L_{h}L_{v}}$, such that $d_{l}=\zeta_{l}e^{-l\rho_{l}}$, such that $|\zeta_{l}|=1$ and $\rho_{l}$ represents the phase shift induced by the $l$-th reflecting element of the IRS $\forall\ 1\leq l\leq L_{h}L_{v}$. It is important to note that the total path delay $\tau_{k,l}>\tau^{\prime}_{k,l}$, and is given by
$\tau_{k,l}=\frac{\|\mathbf{r}_{k}-\mathbf{r_{\mathrm{R}}}\|_{2}+\|\mathbf{r}_{k}-\mathbf{r_{\mathrm{I}}}\|_{2}+\|\mathbf{r_{\mathrm{I}}}-\mathbf{r_{\mathrm{R}}}\|_{2}}{c}.$ (3)
@@
where $\mathbf{v}(t,t_{l})$ is an additive white Gaussian noise (AWGN) vector. For conventional targets, $\alpha^{\mathrm{ms}}_{k}$ is typically large; hence, returns from both paths must, in principle, be processed. However, since the monostatic and IRS-assisted paths incur different delays, the corresponding contributions can be separated and processed independently. This separation can be achieved via beam-space processing. Because the direction (and hence the spatial signature) of the IRS relative to the phased-array radar is known, we construct $M$ receive beams steered toward the IRS with some degree of overlap. The beam-space output sampled at fast-time interval $T_{s}$ is then given by
$\begin{split}\mathbf{\overline{y}}(nT_{s},t_{l})&=\mathbf{W}\big\{\mathbf{y}^{\mathrm{ms}}(nT_{s},t_{l})+\mathbf{y}^{\mathrm{is}}(t,t_{l})+\mathbf{v}(t,t_{l})\big\}\\ &\approx\sum_{k=1}^{K}\alpha^{\mathrm{is}}_{k}\mathbf{\overline{F}}\mathbf{D}\mathbf{h}_{k}\mathbf{g}_{k}^{H}\mathbf{s}(nT_{s}-\tau_{k,l})+\mathbf{\overline{v}}(nT_{s},t_{l}),\end{split}$ (5)
where $\mathbf{\overline{F}}=\mathbf{W}\mathbf{F}$ and $\mathbf{W}\in\mathbb{C}^{M\times M_{R}}$ is the beam-space matrix whose beamforming vectors preserve (primarily) the signal arriving from the IRS direction. We assume that the monostatic component is sufficiently suppressed by these IRS-pointing beams.
Since parameter extraction from conventional targets using the monostatic path is well studied, we focus in this paper on target parameter estimation using the IRS-path signal in (5). The radar data cube $\mathbf{Y}\in\mathbb{C}^{M\times N\times L}$, where $N$ and $L$ denote the numbers of fast-time and slow-time samples, respectively, is processed via a discrete Fourier transform (DFT) along the slow-time dimension to obtain $M$ Range-Doppler (RD) maps, $\mathbf{Y}_{\mathrm{RD}}=\mathcal{F}_{L}\{\mathbf{Y}\}$, where $\mathcal{F}_{L}\{.\}$ is an $L$-point DFT operator. Due to the superposition of the $K$ target responses, $\mathbf{Y}_{\mathrm{RD}}=\sum\limits_{k=1}^{K}\mathbf{Y}^{k}_{\mathrm{RD}}$, where $\mathbf{Y}^{k}_{\mathrm{RD}}$ denotes the RD contribution due to the $k$th target alone. A constant false-alarm-rate (CFAR) [9] detector (or its variants) provides $K_{P}\leq K$ RD bins corresponding to true detections; some bins may contain multiple targets. Therefore, $K=\sum_{p=1}^{K_{P}}P_{k}$, where $P_{k}$ denotes the number of targets present in the $k$th detected RD bin, $(\hat{R}_{k},\hat{v}_{k})$. Note that $\hat{R}_{k}=\frac{c\tau_{k,l}|_{l=0}}{2}$ corresponds to the round-trip distance via the IRS, while $\hat{v}_{k}$ denotes the relative velocity of the target with respect to the radar.
The observation vector for the $k$th RD bin across the $M$ beams is given by
$\mathbf{\overline{y}}_{k}=\mathbf{\overline{F}D}\sum_{p=1}^{P_{k}}\mathbf{h}_{p}\mathbf{g}_{p}^{H}\mathbf{s}+\mathbf{\bar{v}}_{k},$ (6)
We denote the channel between the radar and the $p$th target as $\mathbf{g}_{p}=\alpha_{p}^{(g)}\mathbf{a}_{M_{T}}(\theta^{(\mathrm{az})}_{p},\theta^{(\mathrm{el})}_{p})$, where $\mathbf{a}_{M_{T}}(\theta^{(\mathrm{az})}_{p},\theta^{(\mathrm{el})}_{p})$ is the steering vector toward the target as seen from the transmit array. Similarly, we write $\mathbf{h}_{p}=\alpha_{p}^{(h)}\mathbf{a}_{L_{h}L_{v}}\left(\psi^{(\mathrm{az})}_{p},\psi^{(\mathrm{el})}_{p}\right)$, where $\mathbf{a}_{L_{h}L_{v}}\left(\psi^{(\mathrm{az})},\psi^{(\mathrm{el})}\right)=\mathbf{a}_{L_{v}}\left(\psi^{(\mathrm{az})}\right)\otimes\mathbf{a}_{L_{h}}\left(\psi^{(\mathrm{el})}\right)$ is the steering vector toward the target as observed from the IRS panel, and $\otimes$ denotes the Kronecker product. Defining $\alpha_{p}=\alpha_{p}^{(h)}\alpha_{p}^{(g)}$, (6) can be rewritten as
$\mathbf{\overline{y}}_{k}=\mathbf{\bar{F}D}\sum_{p=1}^{P_{k}}\alpha_{p}\mathbf{a}_{L_{h}L_{v}}\left(\psi_{p}^{\left(\mathrm{az}\right)},\psi^{\left(\mathrm{el}\right)}_{p}\right)\\ \times\mathbf{a}_{M_{T}}\left(\theta^{\left(\mathrm{az}\right)}_{p},\theta^{\left(\mathrm{el}\right)}_{p}\right)^{H}\mathbf{s}+\mathbf{\bar{v}}_{k}.$ (7)
With $\hat{R}_{k}$ and $\hat{v}_{k}$ obtained from the constant false-alarm-rate (CFAR) detector, the parameters $\theta^{\left(\mathrm{az}\right)}_{k},\theta^{\left(\mathrm{el}\right)}_{k},\psi^{\left(\mathrm{az}\right)}_{k},\psi^{\left(\mathrm{el}\right)}_{k},\forall\ k\in[1,K]$ are to be estimated. Importantly, the 3-D target position can be determined using only a subset of these parameters. Let $R_{k,\mathrm{R}}$ and $R_{k,\mathrm{I}}$ denote the distances from the $k$th target to the radar and to the IRS, respectively. Then the estimated range inferred from the path delay satisfies
$\hat{R}_{k}=R_{k,\mathrm{R}}+R_{k,\mathrm{I}}+R_{\mathrm{I},\mathrm{R}},$ (8)
where the IRS–radar distance $R_{\mathrm{I},\mathrm{R}}$ is assumed known.
It follows that estimating $\psi^{(\mathrm{az})}_{k}$ and $\psi^{(\mathrm{el})}_{k}$, together with $R_{k,\mathrm{I}}$, yields the $k$th target position $\mathbf{r}_{k}^{\prime}=[x_{k}^{\prime},y_{k}^{\prime},z_{k}^{\prime}]$ with respect to the IRS coordinate system. The target position in the radar coordinate system can then be obtained via
$\mathbf{r}_{k}=\mathbf{r}_{k}^{\prime}+\mathbf{r}_{\mathrm{I}}.$ (9)
Alternatively, the target position can be obtained directly by estimating $\theta^{\left(az\right)}_{k}$ and $\theta^{\left(el\right)}_{k}$ along with $R_{k,\mathrm{R}}$ from $\hat{R}_{k}$. In the next section, we present a technique to estimate the target position with respect to the IRS.

### v2

# Cramer-Rao Bounds for Target Parameter Estimation in a Bi-Static IRS-Assisted Radar Configuration

@@ Abstract
The use of Intelligent Reflective Surfaces (IRS) to assist communication and sensing has proven cost-effective in challenging scenarios. For sensing, IRS is shown to sense non-line-of-sight (NLOS) and stealth targets, albeit with significant loss due to the four-hop path model. Amongst the available IRS-assisted configurations, we consider a three-hop model in which the IRS redirects the scattered target response towards the mono-static radar. With the IRS spatially displaced from the radar, this configuration mimics a bi-static radar. While target detection has been studied in this configuration, parameter estimation has not been investigated to date. To this end, we first develop the signal model for this configuration and derive the CRB for target parameters. The dependence of CRB on system parameters such as SNR, number of snapshots, number of IRS elements and their weights is brought forward through extensive simulations. This study can enable a designer to customize the system parameters to meet the requirements. It also serves as a benchmark for parameter estimation techniques developed for this configuration.
@@ I Introduction
Radar has historically been a vital sensing technology capable of active and passive detection, target tracking, high-resolution imaging, and, increasingly, sensing in Non-Line-of-Sight (NLoS) environments [16, 23, 13, 19]. Conventional mono-static radars usually perform well when targets are within their Line-of-Sight (LoS), enabling precise measurement of range, Doppler shifts, and directions. However, when obstacles block the LoS or targets exhibit stealth behaviour, the effectiveness of traditional radar diminishes significantly or results in complete detection failure, driving the innovation of new sensing architectures.
A common method for enabling NLoS radar sensing involves the use of an Intelligent Reflective Surface (IRS). An IRS is made up of passive, adaptable reflecting elements that can steer incident electromagnetic waves, creating controllable pathways beyond the LoS when positioned appropriately. This IRS-assisted radar technology is well suited in defence, automotive sensing, and modern wireless systems to expand the field-of-view (FOV) and improve NLoS target detection capabilities [22, 20, 4]. Meanwhile, the presence of an IRS introduces multiple possible signal paths from the target to the radar. Specifically, deploying an IRS results in four possible propagation paths [17]:
@@
Most IRS-assisted radar studies typically focus on the dominant LoS component (P1) and the IRS-enabled NLoS component (P2) [5]. NLoS sensing via (P2) has been widely studied, including the derivation of Cramér–Rao Lower Bounds (CRLB) for selected target parameters [14]. Despite the beamforming gain yielded by IRS weights in this configuration, it cannot fundamentally mitigate the severe path loss inherent to the four-hop channel, due to the compounded attenuation over the radar–IRS and IRS–target links [12]. This configuration is therefore not suitable for sensing low radar cross-section (RCS) targets and stealth targets.
@@
To mitigate performance degradation due to path loss and scattering from stealth targets, bi-static and multi-static radar configurations are considered. Unlike mono-static radar, in which the transmitter (Tx) and receiver (Rx) are co-located, bi-static and multi-static radars spatially separate the Tx and Rx (s), thereby forming a distributed sensing architecture [10, 18, 6]. This separation enables the receiver to observe the target from a viewpoint that is different from the LoS path between the transmitter and target. In multi-static reception, multiple spatially separated receivers observe the target returns, which can improve detection probability and tracking performance due to spatial diversity [21, 1]. Such configurations are particularly useful for detecting low-observable (stealth) targets and for sensing space objects, making them attractive for modern defence and space applications [15, 7]. Despite these advantages, they introduce important practical limitations, most notably stringent time and phase synchronization requirements between the Tx and distributed Rx (s), along with increased hardware cost and maintenance overhead compared to the mono-static radar.
@@
Upon investigating these trade-offs, we observe the need for a configuration that (i) reduces the effective path loss relative to the double-IRS (P2) route, and (ii) avoids inter-node time synchronization errors inherent in bi-static/multi-static radars. In this work, we focus on the third path (P3), i.e., Radar $\rightarrow$ Target $\rightarrow$ IRS $\rightarrow$ Radar, which is often overlooked but is extremely useful. As will be shown, this path is analogous to a bi-static/multi-static observation geometry because the IRS provides a spatially displaced “listening” viewpoint relative to the radar, yet it remains a passive structure. This motivates a bi-static-like configuration based on IRS-assisted radar. This configuration has been recently presented in [3] for stealth target detection. However, target-parameter estimation within this IRS-assisted framework (P3) has not been studied in the literature.
@@
In this work, the IRS-assisted radar configuration is inspired by the bi-static radar configuration. The position of the IRS panel will coincide with the position of the bi-static receiver which is spatially away from the mono-static radar transceiver. This position is determined by the desired FOV, which is the intersection of the transmitter and the IRS beam patterns. In addition to this requirement, this configuration requires a line-of-sight path between the IRS and the radar transceiver for the IRS to redirect the target’s response towards the radar. A dedicated receive chain can process the signal received from this channel. The sensing chain therefore leverages (P3), mimicking certain advantages of bi-static/multi-static radar while maintaining a mono-static radar front-end, thereby eliminating explicit time synchronization errors between Tx and Rx. Moreover, employing an IRS panel to realize a bi-static-like sensing geometry in challenging scenarios (e.g., stealth target detection) introduces a novel paradigm: from a geometric sensing standpoint, the IRS-assisted propagation path can be interpreted as effectively translating the sensing aperture to the IRS location and subsequently mapping the estimated target coordinates back into the original radar reference frame.
@@
  • The IRS-assisted radar configuration is presented in comparison to bi-static/multi-static radar and existing IRS-assisted radars. The associated signal model is systematically developed and reformulated into a parametrized model suitable for the Cramér-Rao Bound(CRB) derivation.
  • We derive CRB for the target angle parameters (azimuth and elevation) under the proposed signal model.
  • Extensive simulation studies are performed to validate the performance of the IRS-assisted radar configuration under different geometric and channel conditions.
The paper is organized as follows: Section II explains in detail the identified IRS-assisted radar configuration. Section III provides the detailed mathematical derivation of the signal model and the target parameters. Based on these target parameters, the theoretical CRB limits are derived in Section IV. Section V discusses the numerical results in relation to the theoretical bounds, and Section VI concludes the paper.
@@ II IRS-Assisted Radar Setup
Bi-static radar separates the transmitter and receiver, forming a signal path with the target as shown in Fig. 1(a)(i), unlike mono-static radar’s setup . This concept generalizes to multi-static radar in Fig. 1(a)(iii), which uses multiple spatially dispersed transmitters and receivers to gain spatial diversity. Multi-static systems observe targets from several angles, enhancing the detection of complex or stealth targets whose Radar Cross Section (RCS) varies with aspect angle. However, effectively exploiting these advantages requires careful spatial arrangement of the receivers, as discussed in [9].
The analogous IRS-assisted radar setup replaces traditional bi-/multi-static receiver nodes with low-cost IRS panels, as shown in Fig. 1(a)(ii) and Fig. 1(a)(iv) that redirects the scattered signals toward nearby radar transceivers [3]. By carefully designing the IRS element weights, the system achieves sufficient beamforming gain to offset longer propagation paths. As shown in Section V, the parameter estimation performance depends on the number of IRS elements with appropriate weights. Fixed IRS placement simplifies radar-IRS channel estimation, and facilitates circumventing synchronization issues typically observed in conventional bi-/multi-static radars.
Amongst the four IRS-assisted configurations discussed in [2] ((P1)-(P4)), the considered IRS-assisted configuration caters to (P1) and (P3) signal paths. This enables the detection of both conventional and stealth targets. Unlike prior IRS-assisted radar work [8], which places the IRS near the radar or target to boost signal strength and relies on a 4-path model (P2) [14], the proposed design positions the IRS panel away from the radar and operates it in simplex mode so that the target response follows the Target → IRS → Radar path. This yields an effective three-path model (P3), reduces path loss compared to conventional four-path IRS-assisted systems, and functionally emulates a bi-static radar configuration. We therefore address this configuration of interest as “IRS-static radar” in the rest of the paper.
@@
((c))
Figure 1: Comparison of radar configurations (left), IRS-static radar setup (center), and IRS-static radar field-of-view (FOV) (right).
@@ III Signal Model
Consider a scenario with $K$ targets, with the radar located at $\mathrm{\textbf{r}}_{\mathrm{R}}=[0,0,0]^{T}$ of the coordinate system. Let the $k$th target parameters-namely range, azimuth angle, and elevation angle - be given by $\textbf{r}_{k}$, $\theta^{(\mathrm{az})}_{k}$, and $\theta^{(\mathrm{el})}_{k}$, respectively. Without loss of generality, we consider a 2-D IRS panel, with $L=L_{h}L_{v}$ elements, where $L_{h}$ and $L_{v}$ denote the number of elements along the horizontal and vertical dimensions, respectively, located at $\mathbf{r}_{\mathrm{I}}$ in the radar coordinate system. Let the $k$th target subtend azimuth and elevation angles $\psi^{(\mathrm{az})}_{k}$ and $\psi^{(\mathrm{el})}_{k}$, respectively, as seen from the IRS coordinate system, illustrated in Fig. 1(b). We assume a phased-array radar with $M_{T}=M_{R}$ transceivers co-located at $\mathbf{r}_{\mathrm{R}}$.
@@
Next, we define the channel between the $k$th target and the IRS panel as $\mathbf{h}_{k}\in\mathbb{C}^{L_{h}L_{v}\times 1}$, and the channel between the IRS panel and the phased-array transceiver as $\mathbf{F}\in\mathbb{C}^{M_{R}\times L}$. The signal received at the radar via the IRS path (denoted as “is”) is given by
$\mathbf{y}^{\mathrm{is}}(t,t_{l})=\sum_{k=1}^{K}\alpha^{\mathrm{is}}_{k}\mathbf{F}\mathbf{D}\mathbf{h}_{k}{\mathbf{g}_{k}}^{H}\mathbf{s}(t-\tau_{k,l}).$ (2)
Let $\alpha^{\mathrm{is}}_{k}$ denote the response strength associated with the IRS-reflected path and the diagonal matrix $\mathbf{D}=\mathrm{diag}(d_{1},\dots,d_{l},\dots,d_{L_{h}L_{v}})\in\mathbb{C}^{L_{h}L_{v}\times L_{h}L_{v}}$ contains the reflection coefficients of all IRS elements. Each coefficient is expressed as $d_{l}=\beta_{l}e^{-j\rho_{l}},$ where $|\beta_{l}|=1$ denotes the unit-modulus amplitude, and $\rho_{l}$ is the phase shift induced by the $l$-th reflecting element of the IRS, for all $1\leq l\leq L_{h}L_{v}$. It is important to note that the total path delay $\tau_{k,l}>\tau^{\prime}_{k,l}$, and is given by
$\tau_{k,l}=\frac{\|\mathbf{r}_{k}-\mathbf{r_{\mathrm{R}}}\|_{2}+\|\mathbf{r}_{k}-\mathbf{r_{\mathrm{I}}}\|_{2}+\|\mathbf{r_{\mathrm{R}}}-\mathbf{r_{\mathrm{I}}}\|_{2}}{c}.$ (3)
@@
where $\mathbf{v}(t,t_{l})$ is an additive white Gaussian noise (AWGN) vector. For conventional targets, $\alpha^{\mathrm{ms}}_{k}$ is typically large; hence, returns from both paths must, in principle, be processed. However, because the path delays differ between the monostatic path and the path via the IRS, the signals must be processed independently. This can be accomplished by beam-space processing or by having a dedicated receive array looking towards the IRS panel. Since parameter extraction from conventional targets using the monostatic path is well studied, we focus our study on the parameter estimation from the IRS path. The signal received by this array from only the IRS path is given by:
$\small\begin{split}\mathbf{\overline{y}}(nT_{s},t_{l})&=\sum_{k=1}^{K}\alpha^{\mathrm{is}}_{k}\mathbf{F}\mathbf{D}\mathbf{h}_{k}\mathbf{g}_{k}^{H}\mathbf{s}(nT_{s}-\tau_{k,l})+\mathbf{v}(nT_{s},t_{l}).\end{split}$ (5)
It is observed that the IRS weights in $\mathbf{D}$ influence the received signal in (5). We next investigate the design of the IRS coefficients and the reconstruction of the underlying signal model in a manner that facilitates estimation of the target parameters in the following subsections.